### v1

# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

Abstract
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet it is challenging to determine what governance strategies are most successful, as measuring the ‘success’ of governance is complex and nuanced. Furthermore, the incredible diversity in community topic, size, and membership all but guarantee that there is no ‘one-size-fits-all’ solution for community governance. In this work, we measure governance by assessing how community members publicly discuss their own moderators. We quantify perceptions of moderators through 1.89 million labeled posts and comments made on reddit over an 18 month period, and relate these perceptions to characteristics of community governance and to different actions that community moderators can take. We identify key differences between different types of communities, and highlight promising strategies for moderator teams. Amongst other findings, we show that positive perceptions of moderators are associated with other measures of community health, and that strict rule enforcement is perceived more favorably for certain topics, such as news communities, than others. We investigate what kinds of moderators have the most positive impact on the community when they join the mod team, and find that moderators who are active community members before and during their mod tenures result in the largest improvement of community members’ perceptions of moderators. We make all our models, datasets, and code public.
1 Introduction
Millions of online communities rely on volunteer moderators to carry out governance. These moderators perform essential services that substantially shape their communities, including setting and enforcing rules (Fiesler et al. 2018; Chandrasekharan et al. 2018; Seering et al. 2019), soliciting feedback from and communicating their actions to the community (Jhaver, Bruckman, and Gilbert 2019; Zhang, Hugh, and Bernstein 2020), and recruiting new moderators to keep pace with community growth or to replace moderators who have stepped down. Some form of governance provided by moderators seems essential to the healthy functioning of almost all online communities (Kraut and Resnick 2012; Matias 2019a).
It is challenging to determine what styles of governance are most effective at ensuring high quality outcomes for online communities. The solution space for online governance is extraordinarily large, with moderator teams needing to make countless decisions about how to govern: how many moderators do they need, and how should they select them? What rules should they enforce, and how should they enforce them? What even is the community’s best interest, anyways? Compounding this challenge is the enormous scale and diversity of online communities, millions of which exist on many platforms, in many sizes, and with many topics and goals. Prior work has established that there is no ‘one-size-fits-all’ solution for improving online governance (Weld, Zhang, and Althoff 2024, 2022).
This scale and diversity also presents an opportunity. If we can develop an effective method to scalably assess the success of community governance practices, we can leverage the natural diversity of these practices across thousands of communities in order to identify the most promising strategies for moderators. However, measuring the success of governance practices at scale is extremely difficult. Survey methods can be effective but expensive to deploy widely, challenging to run longitudinally, and impossible to ‘backfill’ historic data. Notions of ‘success’ are multifaceted, and while automated classifiers can be used to detect specific harms such as misinformation (Weld, Glenski, and Althoff 2021) or specific aspects of governance such as rules enforcement (Chandrasekharan et al. 2018; Fiesler et al. 2018), current methods are unable to quantify broader notions of successful governance.
In this work, we measure online community governance by examining how community members themselves perceive their moderators. We develop a method to classify how community members discuss their moderators, publicly, within their communities (§3). We use this method to gather and characterize community members’ perceptions of moderators at a massive scale, enabling the largest study-to-date of governance practices that we are aware of. We label 1.89 million posts and comments from 8,477 unique subreddits across an 18-month period from January 2021-June 2021, and relate these data to different kinds of online communities, and to different actions that community mod teams can take.
Our analyses address two key research questions:
  • RQ1
    How are moderators of different communities perceived differently by their communities? (§4)
  • RQ2
    What can moderators do to improve how they are perceived by their community? (§5)
We find that community members’ perceptions of their mods vary substantially from community to community (§4). Hobby communities have the most positive perceptions of their mods, with meme and news communities having the poorest. Communities that perceive themselves as having high quality content, and being trustworthy, engaged, inclusive, and safe all perceive their mod teams more positively than communities at consider themselves low quality, untrustworthy, disengaged, uninclusive, or unsafe. Community size is also a major differentiator—tiny communities (1-10 posts+comments/day) use 6.1×6.1$\times$6.1 × as much positive language to describe their mods as huge communities (¿10k posts+comments/day).
Our findings have important implications for moderator teams (§5). We show that communities with fewer than 5 daily posts+comments per moderator use 2.5×2.5$\times$2.5 × as much positive language to describe their mods as communities with 20×20$\times$20 × more posts+comments per mod. Yet our findings do not suggest that more strict moderation improves perception of moderators; in fact, for most types of communities, a 3 percentage point increase in removed content is associated with a 33% increase in negative language used to describe mods, although news-sharing communities are a notable exception to this trend (§5.1). We examine the impact of adding different types of moderators to a community, and find that moderators who are also active community members before and during their tenure as moderators are most strongly associated with improved perceptions of mods (§5.2-9).
We discuss the limitations of our methods, and identify key areas for future work (§6). We make our models, datasets, and code public [^1] These materials are being prepared for publication. For immediate use, please contact the first author. to enable further research on this important topic.
2 Related Work
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera, Halfaker, and Terveen 2009) and some gaming servers (Frey and Sumner 2019). On reddit and most popular social media platforms, however access is difficult to obtain, and, and requires obtaining authorization separately from the mods of every community to be studied (Li, Hecht, and Chancellor 2022a, b). In contrast, our method works for every community whose content is public.
Because of these challenges, many researchers turn to surveys to study governance, sometimes qualitatively (Matias 2019a), and sometimes at a larger scale, for instance to quantify harassment of mods (Almerekhi, Kwak, and Jansen 2020) or moderator recruiting practices (Seering and Kairam 2022).
Governance Through the Lens of Rules. Partially because of their more-public nature, rules are often used as a lens to study governance. Large-scale analyses of rules have been used to characterize governance at the platform level on reddit (Reddy and Chandrasekharan 2023; Fiesler et al. 2018) and on Wikipedia (Hwang and Shaw 2022), however it is challenging to infer much about governance in specific communities, as rules and their enforcement constitute only a small portion of governance activity. Surveys of community members’ attitudes towards rules have been used to understand members’ attitudes towards governance more broadly (Koshy et al. 2023), however this method is challenging to scale beyond a single community. Our method can easily be applied to thousands of communities at once.
Moderation Strategies on Reddit. Rules and their enforcement can also be evaluated in terms of their impact on the community as a whole (Jhaver, Bruckman, and Gilbert 2019; Srinivasan et al. 2019; Jhaver, Rathi, and Saha 2023) and their embodiment of communitywide norms (Chandrasekharan et al. 2018). Beyond rules and their enforcement, researchers have also studied other moderator strategies on reddit, including platform level decisions such as community bans and quarantines (Chandrasekharan et al. 2020; Ribeiro et al. 2020), and community level actions such as user bans (Thomas et al. 2021) and stickied posts (Matias 2019b). Our work quantifies many of these strategies, including content removal, mod interactions with the community, and associates them with positive and negative moderation discourse.
Measuring Outcomes in Online Communities. Researchers have studied the range of values that community members hold for their communities (Weld, Zhang, and Althoff 2024), yet it is challenging to accurately predict these values automatically (Weld, Zhang, and Althoff 2022). Large-scale embeddings can be used to understand community culture (Waller and Anderson 2020), yet not governance directly. Longitudinal work has examined how communities fare with massive growth (Lin et al. 2017) and the lifecycles of their members(Danescu-Niculescu-Mizil et al. 2013), while some research focuses on smaller scale outcomes at the conversation level (Bao et al. 2021). Our work complements this literature with a large scale method for studying governance-specific outcomes.
3 Methods
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al. 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, and moderators. These attributes, along with the the fact that almost all content on reddit is publicly available, make it an ideal platform for studying community governance at a large scale.
In this work we make extensive use of publicly available reddit data downloaded from the Pushshift dataset (Baumgartner et al. 2020). From these data, we detect and classify and detect posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). Additionally, we quantify moderators’ activity on the platform before and during their moderator tenures (such as the number of times a mod posted in a community before being appointed a moderator there, as well as activity by non-moderators in thousands of communities).
3.1 Computing Moderator Timelines
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of each subreddit’s moderators using the Wayback Machine, a web archiving service provided by the Internet Archive. [^2] https://archive.org/ We start by scraping current moderator info pages for the 10,000 largest subreddits on the platform, directly from reddit. Each such page contains a list of the current moderators, in order of seniority, along with the exact timestamp that each moderator was added as a moderator to the subreddit. We then use the publicly accessible API from the Internet Archive’s Wayback Machine to scrape every archived copy of every subreddits’ moderator info page going back to 2010. We scrape 30,302 historical archived copies of moderator info pages, which, combined with the 10,000 present-day copies (one per subreddit), give us detailed timelines of who moderated each subreddit when, and when they started as a moderator. Due to the functionality of the Wayback Machine, we have higher temporal resolution (more archived snapshots) for larger and more popular subreddits, yet, because each moderator info page encodes the exact timestamp for each moderator’s start date, having fewer snapshots results only decreased accuracy of when moderators resign their posts, not when they are appointed — which is the primary focus of this paper. Since the exact end time of a moderator’s tenure can only be inferred from examining when they were removed from the list of moderators for each subreddit, we adopt a conservative strategy which deliberately underestimates the length of moderators’ tenures: we consider the end date of their tenure as that of the last snapshot for which they were still listed as a moderator. Our method only misses moderators whose appointment, entire tenure, and resignation all occur within the ‘gap’ between snapshots [^3] We believe this is rare, as the average mod tenure we can detect is longer than four years, and the mean gap between snapshots is 56 days.. After a final data-cleaning step to remove known bots, the resulting dataset of moderator timelines, which we make public, consists of 89,237 moderator tenures across 9,502 subreddits.
Human Random VADER GPT-4 Our Model
Positive $F_{1}$ 0.85 0.14 0.30 0.70 0.71
Neutral $F_{1}$ 0.89 0.50 0.34 0.71 0.73
Negative $F_{1}$ 0.84 0.13 0.34 0.61 0.71
Exclude $F_{1}$ 0.98 0.21 0.00 0.66 0.73
Test Set Acc. 89.0% 27.0% 30.6% 66.2% 72.4%
Table 1: Our Classification Step model, a LLaMA 2 model fine tuned with QLoRA (Touvron et al. 2023; Dettmers et al. 2023), exceeds the performance of a retrieval-based few-shot classifier using GPT-4. Our model, which we make public, is alsomore scalable (as it can be deployed locally), more affordable, and more reliable (as it is not subject to prompt filtering), than GPT-4. This table compares $F_{1}$ scores for expert human labelers, retrieval-based few-shot GPT-4, and our model, alongside an empirical class distribution random baseline and a VADER-based classifier (Hutto and Gilbert 2014).
3.2 Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the 10,000 largest subreddits, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld, Zhang, and Althoff 2022). We used the manually-labeled dataset of 123 subreddits from Weld, Zhang, and Althoff (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld, Zhang, and Althoff 2022). We average across all responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
3.3 Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify posts and comments made by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects posts and comments which discuss moderators (differentiating them from those which discuss other uses of ‘mod(s)’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of these posts and comments into positive, neutral, negative, and exclude classes. We only include comments made by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not on how mods discuss themselves in public. Our pipeline, which we make public, was applied to the entire reddit dataset of public posts and comments made from January 2020 to June 2021.
Figure 1: Determining the sentiment with regards to the moderators of comments can be very challenging.
Detection Step Details (Step 2). To differentiate posts and comments which discuss moderators from those which use ‘mod(s),’ we fine-tuned a RoBERTa-based binary classifier (Liu et al. 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the entire reddit dataset and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88, and, when applied to the entire reddit dataset, left us with 330,356 posts and 2,426,049 comments which discuss moderators.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa. To manually label posts and comments for the sentiment classification task, two annotators worked together to iteratively refine a codebook (Appendix A), then independently labeled 200 randomly sampled posts and comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts and comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts and comments for use as a training set, of which 484 were randomly sampled, and additional 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa based model trained for the classification task.
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the appropriate moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators (e.g., discussion of the /r/cats moderators taking place on /r/cats). For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with posts or comments that erroneously slipped through the detection step. We trained on this ‘exclude’ class in addition to the classes used for downstream analyses: positive, neutral, and negative [^4] Additionally, we attempted to classify specific complaints about moderation: excessive moderation, insufficient moderation, and biased moderation, but found classifier performance not to be amenable for high-confidence downstream analyses..
To further enhance the performance of our model, we perform data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al. 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al. 2023) fine-tuned with QLoRA (Dettmers et al. 2023). LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step leave us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make this dataset public.
Accurately classifying the sentiment with regards to the moderators of a post or comment is an extremely challenging task that is often heavily reliant on context, background knowledge of the community in which the discussion is occurring, and requiring inside knowledge about the community’s norms and culture (Figure 1). As such, off-the-shelf sentiment classifiers, such as VADER, perform very poorly (Table 1). Even general performance LLMs such as GPT-4, which obtain state-of-the-art results on standard sentiment analysis benchmarks (Kheiri and Karimi 2023), perform worse than our fine-tuned LLaMA2 QLoRA model, even when prompted using a retrieval-based few-shot in-context learning method.
Figure 2: Communities that consider themselves higher quality (a), more trustworthy (b), more engaged (c), more inclusive (d), and more safe (e) all use more positive and less negative sentiment to describe their moderators. Here, communities are grouped into quartiles based on their community members’ self-reported perceptions of the current state of the community. This effect is most pronounced for communities’ self-reported trustworthiness (b), with the top-25% most trustworthy communities using 34% more positive and 22% less negative language to describe their mods. Communities that rate themselves as feeling smaller (f) have a more positive perception of their mods.
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4&4: The Amount of Mod Discourse for a given community and a given time period is the fraction of all posts and comments within that community made during that time period which are discussing mods, regardless of their sentiment. The Fraction of Mod Discourse with positive/negative sentiment is the fraction of all posts and comments discussing mods in a given community that have positive/negative sentiment. For many of our analyses, the time period used is the 18-month period from January 2020 through June 2021. To ensure a meaningful amount of data, we exclude communities that did not have at least one positive, neutral, and negative post or comment discussing moderators. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. These filtering steps leave us with a group of 5,282 subreddits used in downstream analyses. For some analyses, we also compute the Impact of Particular Mod’s Appointment by computing the fraction of mod discourse for the 8-week period immediately preceding and following a new moderator’s appointment to a community, and then taking the difference before vs. after they were appointed.
3.4 Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. Additional dataset details, including FAIR principles, are in Appendix C. We comply with relevant licenses for NLP models we use or modify. This study was approved by the University of Washington IRB under ID number STUDY00011457.
4 How are moderators of different communities perceived differently by their communities?
Online communities exist for nearly every conceivable topic, and range in size from just a few members to many millions. In this section, we examine how community members’ perceptions of their moderators vary across communities with different topics, different community health metrics, and different sizes.
Method. To quantify a community’s health, we divide communities into quartiles based on the responses of their community members to a recent survey assessing the current state of 2,151 communities (Weld, Zhang, and Althoff 2022). Survey responses were collected between May-July 2021, a period overlapping the the end of our mod discourse data time range. For this analysis only, we exclude communities which were not surveyed. Community topic is classified into one of six topical categories using our topic classifier (§3.2). We quantify community size as the volume of submitted content: the average number of posts and comments per day over our study period, which we use to group communities with similar size. For each group, we compute the average fraction of mod discourse, both positive and negative, as well as the total amount of mod discourse.
Results. Different aspects of community health are associated with better and worse perceptions of moderators. The smallest-feeling 25% of communities, based on member self-reports, use 27% more positive and 16% less negative sentiment to discuss their moderators than the largest-feeling 25% of communities (Figure 2a). Figure 2a-e shows that communities which rate themselves as having higher quality content and being more trustworthy, more inclusive, and more safe all use more positive and less negative to discuss their moderators, as well. The difference between the top- and bottom- quartiles are most pronounced for the communities that consider themselves the most- vs. least- trustworthy, with the most-trustworthy communities using 34% more positive and 22% less negative sentiment to describe their moderators (Figure 2b).
Figure 3: Moderators in small communities are perceived more positively and less negatively than moderators in large communities. (a) Tiny communities’ discussion of their moderators has 6.0×6.0$\times$6.0 × more positive sentiment than huge communities, while huge communities further use 2.1×2.1$\times$2.1 × more negative sentiment to describe their mods, relative to the total amount of mod discourse. (b) Tiny communities also tend devote a relatively large proportion of their content to discussing their moderators, with tiny communities having on average 6.1×6.1$\times$6.1 × more posts and comments discussing moderators than larger communities.
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities seem to have more favorable perceptions of their moderators (Figure 3a), with tiny communities with $<10$ posts+comments/day using 6.0×6.0$\times$6.0 × more positive and 0.46×0.46$\times$0.46 × as much negative sentiment to describe their mods than huge communities (those with more than 10k posts and comments per day), relative to the total amount of mod discussion. Tiny communities also devote 7.6×7.6$\times$7.6 × as many total posts and comments to discussing their mods than huge communities, although this effect diminishes with increasing community size—the average amounts of mod discussion in medium, large, and huge communities are all within ±$\pm 7.0\%$ of one another (Figure 3b).
Examining communities of different topics, hobby communities have generally the most positive perception of their moderators, while meme and news-sharing communities have the most negative perception, with meme and news-sharing communities having 46% more of their mod discourse have negative sentiment (Figure 4). Meme and news sharing communities also have more mod discourse than hobby, identity, and photo/video communities. Discussion communities have the most mod discourse of any community topic, on average (Figure 4).
Figure 4: Moderator perceptions varies significantly across communities with different topics. (a) On average, hobby communities have the best perception of their moderators, while meme and news communities have the poorest. (b) Discussion communities have largest amount of mod discourse of any community topic. In this and all other figures, points represent mean estimates alongside bootstrapped 95% confidence intervals.
Implications. Surveys of redditors have shown that they consider a wide range of factors to be important to the overall ‘health’ of their communities (Weld, Zhang, and Althoff 2024). Here we show that communities that are doing well with regards to factors widely considered to be important, such as safety and quality of content, tend to perceive their moderators more positively. Many of these factors are only indirectly controlled by the moderators, for example quality of content—while moderators can set and enforce rules aimed at improve the quality of content in their communities, quality of content is ultimately a function of the content submitted by community members, not moderators. This can lead to moderators being ‘blamed’ for problems largely outside of their control, which has been previously shown in small-n surveys (Matias 2019a; Jhaver, Frey, and Zhang 2023) and is further supported empirically here, as shown in Figure 2.
Small communities appear to both discuss their moderators more, as well as use more positive sentiment in their discussions (Figure 3). Several factors may contribute to this, including that the increased anonymity that comes with participating in a larger community may make people feel more comfortable speaking negatively about the moderators, and that smaller communities are more likely to be newly formed and thus still establishing moderation norms, leading to more mod discourse (Hwang and Foote 2021; Seering and Kairam 2022).
Moderating an online community is a challenging task, with many factors outside the moderators’ control. Some community topics are simply more contentious than others, and therefore more challenging to moderate. We suspect this is a contributing factor to apparent differences in perception of moderators between communities with different topics (Figure 4). To control for this, in subsequent analyses, we validate our findings by stratifying the analyses by community topic and ensuring that trends in results are qualitatively similar, except when stated otherwise.
5 What can moderators do to improve how they are perceived by their community?
Moderators have a great deal of autonomy to run their communities as they see fit, including enforcing rules by removing content and growing the mod team by recruiting or appointing new moderators. Moderators also may (or may not) participate in the community as ‘regular community members’ in addition to their official mod duties. In this section, we examine the impact that different strategies in each of these topics may have on communities’ perceptions of their moderators, and identify promising suggestions for moderators.
Figure 5: Moderators in communities with lower workloads are perceived more positively and less negatively than moderators in communities with high workloads. Communities with lower moderator workloads (more moderators relative to the amount of content submitted) tend to have more more positive sentiment in their discussion of the moderators, and less negative sentiment. Communities with fewer than five posts and comments per mod per day use 2.5×2.5$\times$2.5 × as much positive sentiment in their mod discourse compared to communities with more than 100 posts and comments per mod per day.
5.1 Content Removal and Moderator Workload
Method. We can identify content removed by moderators in each community by counting the occurrences of ‘[removed]’ posts and comments within each community. By dividing these counts by the total amount of content submitted to that community, we can compute the total percentage of content removed by mods. Using our moderator timelines (§3.1), we can compute the total number of moderator-tenures in any given subreddit over the course of our analysis time period. We also estimate the expected ‘workload’ of each mod by dividing the total daily amount of content submitted to each community by the number of mods available to review that content; while in reality it is unlikely that all moderators share the work of reviewing content evenly (Li, Hecht, and Chancellor 2022a); this metric nonetheless helps us understand the ratio of moderators to content within each community. If workload was shared unevenly, this would only make the workload even higher for a subset of moderators, leading to at worst to overly conservative estimates (higher workload to be addressed by fewer moderators).
Results. We find that communities with higher mod workloads are associated with a decreasing fraction of negative mod discourse (Figure 5). Communities with fewer than five pieces of daily content for each mod (8.8% of all communities) have more mod discourse with positive sentiment than with negative sentiment, a rarity on a platform where mods are far more commonly discussed negatively. As the mod workload increases and the number of mods decreases relative to the amount of content in the community, the fraction of positive mod discourse decreases significantly, with communities with more than 100 posts and comments per mod per day using positive sentiment to describe their mods only 0.39×0.39$\times$0.39 × as often as the 8.8% of communities with the lowest mod workload.
Generally, we find that the fraction of mod discourse with negative sentiment is higher in communities with more removed content (Appendix Figure 10). However, these trends vary depending on the topic of the community (Figure 6). [^5] Similar figures for each of the six community topics are included in Appendix Figure 11. Compared to all other communities, news communities seem to reverse this trend, with perceptions of moderators improving as the amount of content removed from news communities increases. Amongst news communities that remove $<0.5\%$ of their content, 60% of mod discourse has negative sentiment, which drops 35% to 39% for news communities whose mods remove between 1.5% and 3% of their content. For the same amounts of removed content for non-news communities, the fraction of mod discourse with negative sentiment jumps 32%, from 24% to 32%.
Figure 6: For most communities, the fraction of mod discourse with negative sentiment increases as community moderators remove more content (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment actually dropping 35% from news communities whose mods remove $<0.5\%$ of content to communities whose mods remove between 1.5% and 3% of content (b).
Figure 7: Newly appointed mods are associated with a greater improvement in mod perceptions if they are engaged in the community and elsewhere on reddit before their tenure, and if they are engaged during their tenure. (a-b) On average, adding a moderator to a community results in a decrease in the fraction of mod discourse in that community with negative sentiment (red), regardless of the degree to which that moderator has engaged with the community prior to, or will engage with during, their tenure. However, adding a moderator who already has or will engage with the community also results in an increase in the fraction of mod discourse in the community with positive sentiment (aqua), and that increase is 32.5% larger when adding a mod who will engage with the community going forward (b) than for one who already has (a). Adding a moderator who is an active member of communities other than the one they are becoming a mod of results in an increase in positive, and a decrease in negative, sentiment in mod discourse.
Implications. Our results suggest that adding additional moderators to a community (and therefore reducing the effective moderator workload) may improve community members’ perceptions of their mods (Figure 5), although our correlational results may or may not imply causation. In §5.2 we report on additional longitudinal evidence that recruiting additional moderators can have a positive effect. However, simply increasing the amount of content which is removed does not, in general, appear to be associated with more positive moderator discourse sentiment—in fact, the opposite appears to be the case for communities that are not focused on sharing news (Figure 6). Taken together, these results imply that there are topic-specific nuances to content removal, and that mods should use care when deciding to remove more or less content, e.g., by increasing or decreasing the strictness by which rules are enforced. Our results suggest that certain topics are more amenable to stricter rule enforcement than others. While the members of news communities feel less negatively about their mods as the mods remove more content, members of discussion communities whose mods remove $\approx 1.5\%$ of content use 1.75×1.75$\times$1.75 × as much negative-sentiment language to discuss their mods as members of discussion communities whose mods remove $<0.5\%$ of content–a steep increase (Appendix Figure 11).
5.2 Community Engagement
Method. Some moderators are actively engaged with community, soliciting feedback from non-moderators, updating community members on moderation-related news, and contributing to regular community content in addition to their official moderator duties. Other moderators are far less visible, opting to remove content and change rules without participating in the community more broadly. Furthermore, many moderators also participate in other communities beyond just the one(s) they moderate. Using moderators’ public posts and comments, for each moderator-appointment to a community, we compute the number of posts and comments they made in that community and in other communities before and during their tenure as a moderator. We combine these counts with the change in the fraction of mod discourse immediately before and after each mod’s appointment (§3.3) to associate the impact that each mod had on the community to which they were appointed.
Results. We find that, regardless of the moderators’ previous engagement (or lack thereof) with the community, adding a new moderator to a community results in a decrease in the fraction of mod discourse with negative sentiment (Figure 7). However, not all moderators have the same impact. The most positively impactful moderators are those who are engaged with the community prior to their appointment (Figure 7a), will continue to engage with the community during their tenures (Figure 7b), and were also active in other communities prior to their appointment (Figure 7c). New mods who do all three of these things are associated with a 2.5pp increase in the fraction of positive mod discourse, whereas mods who do none of these things are instead associated with a 0.5pp decrease.
Implications. Our results suggest that moderators who actively engage with the community both before and during their tenures are most likely to have a positive impact on the community’s perception of their mod team, which may be because these moderators are more familiar with the community, and are better mods as a result. Another plausible mechanism for this effect is that non-moderator community members value the transparency and accountability that may stem from increased moderator engagement. Our results also suggest that moderators who are active in other subreddits beyond the one(s) they moderate have a more positive impact on the communities they moderate, perhaps because participating in a broader range of communities makes them more effective moderators (Zhu, Kraut, and Kittur 2014). Lastly, our results suggest that of these factors, engagement with the community during the moderator’s tenure has the largest impact.
5.3 Moderator Recruiting
When it’s time to grow the mod team, existing moderators have a wide array of options for who to recruit, and how to recruit them, and little guidance, official or otherwise, for how to select new mods. In practice, moderator recruiting falls into two different strategies: Public Recruiting, where moderators post publicly internally in their own subreddit that they are looking for moderators, and solicit applications, nominations, or hold elections. Sometimes, moderators make use of special external moderator-recruiting subreddits, such as /r/needamod, where they can post ‘job listings’ to prospective applicants. By contrast, Private Recruiting is the recruiting of new moderators in the absence of any public recruiting activity. Moderators recruited privately are either invited to join directly by an existing mod, make an offer to moderate themselves, or are recruited via other backchannel means.
Figure 8: Communities that recruit moderators publicly are 8.78×8.78$\times$8.78 × larger than the average community which recruits only privately, in terms of the community’s daily volume of content (a). Small communities lean towards private recruiting. (b) Compared to private recruiting, recruiting moderators publicly is more polarizing, with an increase in both positive and negative fractions of mod discourse, and a corresponding decrease in neutral sentiment.
Method. We identify instances of public recruiting using a regular expression-based search for posts made by sitting moderators which use ‘recruiting new mods,’ ‘applications open for new mods,’ ‘holding mod elections,’ or similar phrases. We also apply a regular expression-based filter to the complete set of posts from /r/needamod to identify which subreddits are recruiting, restricting the results to only posts made by current moderators of the subreddit that is recruiting. We then take our dataset of moderator timelines (§3.1) and match moderators who were appointed to a community within 8 weeks of a public recruiting post as having been publicly recruited. Moderators appointed without a recent public recruiting post are considered privately recruited.
Results. We find that, while communities of all sizes use public recruiting methods, larger communities are the most likely to make use of public recruiting (Figure 8a). In terms of its daily volume of content, the average community that uses both Internal and External Public Recruiting is 8.78×8.78$\times$8.78 × larger than the average community that only recruits privately. Examining the impact that adding a single moderator recruited privately vs. publicly has to a community in Figure 8b, we find that public recruiting appears to be polarizing, with the fractions of both positive and negative mod discourse increasing by 2.7pp and 3.0pp, respectively, after a publicly recruited mod is added, on average.
Implications. Our findings suggest that public moderator recruiting has the potential be a powerful tool to improve community perceptions of moderators when used carefully, and can be harmful when used without regard to the community’s preferences. A plausible explanation for the increased polarization resulting from public recruiting is that in some circumstances, the public nature of the recruiting exacerbates existing frustrations with mod teams, for example if due process was not followed during moderator elections, or if a perceived-outside was brought in via external recruiting when community members themselves preferred someone with more experience in the community. More work is needed to disentangle confounding factors and more deeply assess the differences between different moderators’ appointments.
5.4 Novice and Experienced Mods
Method. When selecting new moderators, existing moderators may be inclined to favor candidates who already have some moderation experience in other communities. Using Moderator Timelines (§3.1), we can accurately assess how much experience a new moderator has at the time the are appointed, if any.
Figure 9: In contrast to small moderator teams, large teams are much more likely to appoint new moderators who already have moderation experience. $94\%$ of mods recruited to join mod teams with more than 100 mods have at least 2 years of experience, while 74% of mods who join small teams with fewer than 10 mods have no previous experience at all.
Results. We find that large communities are much more likely to recruit moderators who already have moderation experience (Figure 9). While $74\%$ of mods appointed to communities with fewer than 10 mods are first time moderators, $94\%$ of mods appointed by communities with $>100$ mods have more than 2 years of experience.
Implications. Large communities with large moderator teams rarely appoint novice moderators, perhaps due to a perception that a large community is not an appropriate place for new moderators to gain experience. As such, it seems that the most common pattern on reddit is for new moderators to start off moderating small communities, and then work their way up to larger ones. This may contribute to a perception, especially amongst non-moderators, that moderators are motivated to increase their number of appointments and to moderate larger and larger communities, potentially biasing the performance of their duties.
6 Discussion
The massive diversity of online communities offers enormous potential to empirically study how to make online communities better. Any such studies, however, require robust and scalable methods to quantify both the outcomes (which communities are doing ‘better’ than others), as well as the independent variables (the aspects of communities that might make them better). Measuring the success of a community’s moderators is more challenging than many other aspects of making a community ‘better,’ as unlike, say, misinformation, governance is far less visible in a community, and its success is far less well defined. The primary contribution of our work is the use community members’ own perceptions of their moderators, leveraging millions of people’s perceptions of good governance, rather than attempting to define good governance on our own terms.
We are tremendously excited about the potential synergies this line of work enables. While the findings presented in this paper are impactful by themselves, combining our measure of governance with many other important measures of community outcomes (safety, inclusion, and discussion quality, for example), along recent advances in causal inference methodology, will enable studies of how to make communities better that are both more comprehensive and more robust than previously possible.
6.1 Limitations
Our work measures community governance through community members’ public discussion of moderators. While these signals enable insights about different governance strategies, they do not capture every aspect of the success of a community’s governance. What people say publicly does not always reflect their actual beliefs, and even if it did, minimizing community unhappiness is not necessarily the best objective function for community moderators. Differences in community members’ behavior and values also may bias our results, as different community may have different norms around mod discourse, and different community members may feel more or less comfortable expressing their opinions publicly. Additional study is needed to ensure that scalable methods for measuring governance reflect the needs of all community members, not just the noisiest.
Even though our fine-tuned sentiment classifier exceeds the performance of state-of-the-art methods (§3.3, it is not perfect. The large volume of data of data used in our analyses minimizes the impact of any single misclassification by the model, and all of our figures include bootstrapped CIs to better understand the robustness of our findings. Furthermore, systems as complex as online communities have countless confounding factors that can bias analuses such as ours. While we attempt to control for confouders by stratifying results by community size and topic and qualitatively assessing the different strata (§4), future work could make use of more sophisticated methods to control for confounders.
Validating measures of online governance is challenging, as there is minimal ‘ground truth’ to use for assessment. While it exceeds the scope of this work, a large-scale survey of many community members’ perceptions of their moderators could be used to refine and validate future models. Future models could go beyond just positive and negative sentiment to identity biased, overly-strict, or too-permissive moderation, for example.
While our methods for quantifying communities’ practices are sophisticated, they also miss many key aspects of governance. Our moderator timelines (§3.1) are limited in their temporal resolution of the end of moderators’ appointments, especially for smaller communities. Our analyses of moderators’ engagement with communities (§5.2) only consider how much each moderator posts in their communities, not what they post. Future analyses could examine the types of contributions that moderators make. Our analyses of moderator team dynamics are also limited by the lack of publicly available data about which moderators are active; many communities’ mod teams contain mods who do not contribute to the day-to-day governance of the community. Future work should incorporate detailed information about specific moderators’ actions, although data collection is a substantial challenge (Li, Hecht, and Chancellor 2022a).
6.2 Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across 8,477 communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§5.1), community engagement (§5.2), and moderator recruiting practices (§5.3-5.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our model and datasets public to support future research.
Acknowledgements
This research was supported by the Office for Naval Research (#N00014-21-1-2154). NSF grant IIS-1901386, NSF CAREER IIS-2142794, NSF grant CNS-2025022, and the Bill & Melinda Gates Foundation (INV-004841).
Appendix A Moderator Sentiment Codebook
A.1 Positive Sentiment
This label should be used for comments expressing a positive sentiment towards the moderator or moderator team.
Examples. “This subreddit is so lucky to have such a great mod team”
“Make the life of our hard-working mods here easier”
“The mods are always so helpful, but this thread got a bit messy” — this is a tricky judgment call, but I’d say that the overall sentiment is positive with this thread being an exception.
Counterexample. “This subreddit used to be well-run, but in the past year or so the moderation has really gone to shit” — a judgment call similar to earlier, but here I would say negative.
A.2 Negative Sentiment
This label should be used for comments expressing a negative sentiment towards the moderator or moderator team.
Examples. “The mods here suck”
“The mods made a mistake” — everyone makes mistakes, but it’s still better if they don’t.
“I’m so tired of mods not removing crap like this”
A.3 Neutral Sentiment
This label should be used when there isn’t enough context for you to make a judgment about the sentiment of the comment or post, or the sentiment seems neutral.
Examples. “I didn’t delete the post, maybe the mods did?”
“Edit: reworded a slur after getting a warning from the mods”
“Mods please ban this person” — not enough explicitly stated sentiment to know what is meant with certainty.
“Why don’t you go and complain to the mods like you usually do?” — negative sentiment, but not directed towards the moderators.
Appendix B Supplementary Figures
Figure 10: On average, communities with a larger fraction of their content removed by mods tend to have more a smaller fraction of their mod discourse have positive sentiment (a), and a larger fraction with negative sentiment (b). Communities with more content removed by mods also tend to have more total mod discourse (c).
Figure 11: Generally, the fraction of mod discourse with negative sentiment is higher in communities with more removed content. However, these trends vary depending on the topic of the community.
Figure 12: On average, adding a new moderator to a subreddit results in an increase in the fraction of mod discourse which has positive sentiment, and a corresponding decrease in the fraction that has negative sentiment. However, the magnitude of the impact varies with the size of the community’s moderator team; adding a single new mod to a community with fewer than 4 mods has a much larger impact than adding a single mod to a community with more mods. Adding a single moderator to a community with more than 15 mods has an impact which is not significantly different from 0.
Appendix C FAIR Principles and Dataset Datasheet
C.1 Compliance with FAIR Principles
Our datasets, which we will make public upon publication of this work, comply with the FAIR Principles (FORCE11 2020).
Findable. Our dataset of mod discourse makes use of existing reddit post and comment ID identifiers. We will publish them with metadata that clearly and accurately describes the data and species the data identifier on a searchable hosting platform.
Accessible. Our datasets are of small enough size ($\approx 5GB$ total) that they can be feasibly downloaded in their entirety using simple and free protocols such as https. No special authentication will be needed to access these data.
Interoperable. Our metadata will be formal, accessible, and applicable, and will use FAIR compliant vocabularies and references.
Reusable. Our datasets are widely usable, with clearly documented attributes, and will be published with an open and clearly stated data usage license.
C.2 Dataset Datasheet
Motivation. These datasets were created to understand community governance on reddit at a large scale.
Composition. Instances composing the mod discourse dataset are posts and comments discussing the moderators of the subreddit each post and comment was made in. Instances composing the moderator timelines dataset represent a moderator’s tenure in a specific subreddit. The mod discourse dataset contains 1.89 million posts and comments, while the moderator timelines dataset contains 89,237 moderator tenures. Each dataset contains all possible instances for their respective time ranges (January 2020 to June 2021 for the discourse dataset, 2006-May 10, 2022 for the timelines dataset). Each instance of the mod discourse dataset consists of a reddit-issued identifier for the associated post and comment, and a ‘predicted sentiment’ label describing the sentiment the post or comment holds towards the moderators: either positive, neutral, or negative. Instances of the mod timelines dataset consist of a moderator’s username, the subreddit they were appointed to, the exact timestamp their appointment started, and an estimated lower bound on the time their appointment ended, based on the archived snapshots the timelines were reconstructed from. To the best of our knowledge, no information is missing. The mod discourse dataset was created using a language model to classify the sentiment of posts and comments with regards to the moderators and thus may contain classification errors. Details of model performance are included in §3.3. The datasets are self contained, although they may be optionally linked to other datasets of reddit content by moderator username and/or post/comment ID. The datasets do not contain confidential or offensive information.
Collection Process. The mod discourse dataset was created by applying a labeling pipeline to raw reddit posts and comments, as described in detail in §3.3. Raw reddit data was obtained using the Pushift dataset dumps (Baumgartner et al. 2020). No sampling was performed in the production of the dataset, and no laborers were used for labeling tasks beyond the authors’ labeling of training and test sets. The timeframe of the data spans all reddit posts and comments made between December 31, 2019, and August 1, 2021. This work was approved by the University of Washington IRB under ID number STUDY00011457.
Preprocessing/Cleaning/Labeling. Several steps of preprocessing were applied to the datasets, as described in detail in §3. While due to licensing and ethical reasons we cannot republish the original raw reddit data used in their production, we make all code used for preprocessing and labeling public.
Uses. Our datasets have already been used for the analyses of reddit governance presented in this paper. A link to this paper will be clearly posted along with the datasets. These datasets can be used for many other analyses of moderation and governance on reddit. We do not believe there are any tasks for which the dataset could reasonably be used that would result in breaches of users privacy or other harms.
Distribution. Our datasets will be distributed publicly upon peer-reviewed publication of this paper. They will be published with an open and clearly stated data usage license. The datasets have no IP restrictions on them, and, as far as we are aware, are not subject to export controls or regulatory restrictions.
Maintenance. Our dataset will be hosted publicly on our University-hosted webpage, as well as on a public academic dataset hosting service to be determined at the time of publication. Contact information for the datset (the corresponding author of this paper) will be made clear and public. Our datasets are derived only from publicly-posted data which we have processed, and we will provide an affordance for users to have their data removed from our dataset.