### v2

# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

Abstract.
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet measuring the ‘success’ of governance is complex and nuanced, making it challenging to determine what governance strategies are most successful. Furthermore, prior work has shown that communities have differing values, suggesting that ‘one-size-fits-all’ approaches to governance are unlikely to serve all communities well. In this work, we assess governance practices on reddit by classifying the sentiment of community members’ public discussion of their own moderators. We label 1.89 million posts and comments made on reddit over an 18 month period. We relate these perceptions to characteristics of community governance, and to different actions that community moderators take. We identify types of communities where moderators are perceived particularly positively and negatively, and highlight promising strategies for moderator teams. Amongst other findings, we show that strict rule enforcement is linked to more favorable perceptions of moderators of communities dedicated to certain topics, such as news communities, than others. We investigate what kinds of moderators are associated with improved community perceptions upon their addition to a mod team, and find that moderators who are active community members before and during their mod tenures are seen more favorably. We make all our models, datasets, and code public.
1. Introduction
Millions of online communities rely on volunteer moderators to carry out governance. These moderators perform essential services that shape their communities, including setting and enforcing rules (Fiesler et al., 2018; Chandrasekharan et al., 2018; Seering et al., 2019), communicating their actions to the community (Jhaver et al., 2019; Zhang et al., 2020), and recruiting new moderators. Some form of governance provided by moderators is essential to the healthy functioning of almost all online communities (Kraut and Resnick, 2012; Matias, 2019a).
Mod teams must determine how many moderators to have, how to recruit new mods, what rules to set, and answer many more questions. With so many decisions to make, it’s challenging to determine what governance practices are most effective at ensuring high quality outcomes for online communities. There are so many different communities, with different sizes and topics, that there is no ‘one-size-fits-all’ solution (Weld et al., 2024, 2022). However, the variety of existing communities presents an opportunity: if we can develop a method to assess the success of communities’ governance practices, we can leverage the natural diversity of these practices to identify the most promising strategies for moderators. Doing so is challenging. Survey methods are effective, but expensive to deploy and cannot enable the use of historic data. Notions of ‘success’ are multifaceted, and while classifiers have been used to detect specific harms such as misinformation (Weld et al., 2021) or specific aspects of governance such as rules enforcement (Chandrasekharan et al., 2018), current methods are unable to quantify broader notions of successful governance.
In this work, we measure online community governance by examining how community members themselves perceive their moderators. We develop a method to classify how community members discuss their moderators, publicly, within their communities (§3). We use this method to gather and characterize community members’ perceptions of moderators at a massive scale, enabling the largest study-to-date of governance practices that we are aware of. We label 1.89 million posts and comments from 8,477 unique subreddits across an 18-month period from January 2020-June 2021, and relate these data to different kinds of online communities, and to different actions that community mod teams can take.
Our analyses address two key research questions:
  • RQ1
    How are moderators of different communities perceived differently by their communities? (§4)
  • RQ2
    What can moderators do to improve how they are perceived by their community? (§6)
We find that community members’ perceptions of their mods vary substantially from community to community (§4). Hobby communities have the most positive perceptions of their mods, with meme and news communities having the poorest. Communities that perceive themselves as having high quality content, and being trustworthy, engaged, inclusive, and safe all perceive their mod teams more positively than communities at consider themselves low quality, untrustworthy, disengaged, uninclusive, or unsafe. Community size is also a major differentiator—tiny communities (1-10 posts+comments/day) use 6.1×6.1$\times$6.1 × as much positive language to describe their mods as huge communities (¿10k posts+comments/day).
We identify actionable strategies for community moderators, using IPTW and DID causal inference methods to control for confounding factors including the topic and size of communities (§5). We show that communities with fewer than 5 daily posts+comments per moderator use 2.5×2.5$\times$2.5 × more positive language to describe their mods as communities with 20×20$\times$20 × more posts+comments per mod. Yet our findings do not suggest that more strict moderation improves perception of moderators; in fact, for most types of communities, a 3 percentage point increase in removed content is associated with a 9pp increase in negative language used to describe mods, although news-sharing communities are a notable exception to this trend (§6.1). We measure the impact of adding different types of moderators to a community, and find that moderators who are also active community members before and during their tenure as moderators are most strongly associated with improved perceptions of mods (§6.2).
We discuss the limitations of our methods, and identify key areas for future work (§7). We make our models, datasets, and code public [^1] Contact the first author for more information. to enable further research on this important topic.
2. Related Work
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera et al., 2009) and some gaming servers (Frey and Sumner, 2019). On reddit and most popular social media platforms, however access is difficult to obtain, and, and requires separate permission from the mods of every community to be studied (Li et al., 2022a, b). In contrast, our method works for every community whose content is public.
Because of these challenges, many researchers turn to surveys to study governance, sometimes qualitatively (Matias, 2019a), and sometimes at a larger scale, for instance to quantify harassment of mods (Almerekhi et al., 2020) or moderator recruiting practices (Seering and Kairam, 2022).
Governance Through the Lens of Rules. Rules are often used as a lens to study governance. Large-scale analyses of rules have been used to characterize governance at the platform level on reddit (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018) and on Wikipedia (Hwang and Shaw, 2022), however it is challenging to infer much about governance in specific communities, as rules constitute only a small portion of governance activity. Surveys of community members’ attitudes towards rules have been used to understand members’ attitudes towards governance more broadly (Koshy et al., 2023), however this method is challenging to scale beyond a single community. Our method can be applied to thousands of communities.
Moderation Strategies on Reddit. Rules and their enforcement can also be evaluated in terms of their impact on the community as a whole (Jhaver et al., 2019; Srinivasan et al., 2019; Jhaver et al., 2023b) and their embodiment of communitywide norms (Chandrasekharan et al., 2018). Beyond rules and their enforcement, researchers have also studied other moderator strategies on reddit, including platform level decisions such as community bans and quarantines (Chandrasekharan et al., 2020; Ribeiro et al., 2020), and community level actions such as user bans (Thomas et al., 2021) and stickied posts (Matias, 2019b). Our work quantifies many of these strategies, including content removal, mod interactions with the community, and associates them with positive and negative moderation discourse.
Measuring Outcomes in Online Communities. Researchers have studied the range of values that community members hold for their communities (Weld et al., 2024), yet it is challenging to accurately predict these values automatically (Weld et al., 2022). Large-scale embeddings can be used to understand community culture (Waller and Anderson, 2020), yet not governance directly. Longitudinal work has examined how communities fare with massive growth (Lin et al., 2017) and the lifecycles of their members(Danescu-Niculescu-Mizil et al., 2013), while some research focuses on smaller scale outcomes at the conversation level (Bao et al., 2021). Our work complements this literature with a large scale method for studying governance-specific outcomes.
3. Methods
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al., 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, membership, and moderators.
In this work we make extensive use of publicly available reddit data from the Pushshift (Baumgartner et al., 2020). From these data, we detect and classify posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). We collect supplementary data about moderators (§3.1) and communities (§3.2).
3.1. Computing Moderator Timelines
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of the 10,000 largest subreddit’s moderators using the Wayback Machine, a web archiving service provided by the Internet Archive (Archive, 2024). We collect archived snapshots of every subreddit’s moderator info page going back to 2010, totalling 30,302 snapshots, which we combine with 10,000 present day snapshots collected directly from reddit. Each snapshot includes a list of moderators’ usernames, ordered by seniority. As the Wayback Machine snapshots higher traffic pages more frequently, we have more higher temporal resolution for larger and more popular subreddits. However, as each moderator info page contains the exact timestamp that each moderator was added, we have second-level precision for new moderator appointments. After a final data-cleaning step to remove known bots, the resulting dataset of moderator timelines consists of 89,237 moderator tenures across 9,502 subreddits.
3.2. Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the 10,000 largest subreddits, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld et al., 2022). We used the manually-labeled dataset of 123 subreddits from Weld et al. (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld et al., 2022). We average across all responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
3.3. Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying posts and comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify content written by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects content discussing moderators (differentiating them from those which make other use of ‘mod’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of this content into positive, neutral, negative, and exclude classes. We only include content written by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not on how mods discuss themselves in public. We apply our pipeline, which we make public, to all reddit posts and comments made from Jan. 2020 to June 2021.
Figure 1. Determining the sentiment with regards to the moderators of comments can be very challenging.
Detection Step Details (Step 2). To differentiate posts and comments which discuss moderators from those which use ‘mod(s),’ we fine-tuned a RoBERTa-based binary classifier (Liu et al., 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the entire reddit dataset and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa (Figure 1). To label posts and comments for the classification task, two annotators worked together to iteratively refine a codebook (Supplementary Materials A), then independently labeled 200 randomly sampled posts/comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch, 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts/comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts/comments for use as a training set, of which 484 were randomly sampled, and 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa model trained for the classification task.
Human Random VADER GPT-4 Our Model
Positive $F_{1}$ 0.85 0.14 0.30 0.70 0.71
Neutral $F_{1}$ 0.89 0.50 0.34 0.71 0.73
Negative $F_{1}$ 0.84 0.13 0.34 0.61 0.71
Exclude $F_{1}$ 0.98 0.21 0.00 0.66 0.73
Test Set Acc. 89.0% 27.0% 30.6% 66.2% 72.4%
Table 1. Our Classification Step model, a LLaMA 2 model fine tuned with QLoRA (Touvron et al., 2023; Dettmers et al., 2023), exceeds the performance of a retrieval-based few-shot classifier using GPT-4. Our model is also more scalable (it can be deployed locally), more affordable, and more reliable (it is not subject to prompt filtering), than GPT-4. This table compares $F_{1}$ scores for expert human labelers, retrieval-based few-shot GPT-4, and our model, alongside an empirical class distribution random baseline and a VADER-based classifier (Hutto and Gilbert, 2014).
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators. For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with content that erroneously passed the detection step. We trained on this ‘exclude’ class in addition to those used for downstream analyses: positive, neutral, and negative.
Figure 2. Communities that consider themselves higher quality (a), more trustworthy (b), more engaged (c), more inclusive (d), and more safe (e) all use more positive and less negative sentiment to describe their moderators. Here, communities are grouped into quartiles based on their community members’ self-reported perceptions of the current state of the community. This effect is most pronounced for communities’ self-reported trustworthiness (b), with the top-25% most trustworthy communities using 34% more positive and 22% less negative language to describe their mods. Communities that rate themselves as feeling smaller (f) have a more positive perception of their mods. In this and all other figures, points represent mean estimates alongside bootstrapped 95% confidence intervals.
To further enhance the performance of our model, we performed data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al., 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al., 2023) fine-tuned with QLoRA (Dettmers et al., 2023). LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step left us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make this dataset public.
Accurately classifying the sentiment with regards to the moderators of a post or comment is an extremely challenging task that is often heavily reliant on context and background knowledge of the community (Figure 1). As such, off-the-shelf sentiment classifiers, such as VADER, perform very poorly (Table 1). Even general purpose LLMs such as GPT-4, which obtain state-of-the-art results on standard sentiment analysis benchmarks (Kheiri and Karimi, 2023), perform worse than our fine-tuned LLaMA2 QLoRA model, even when prompted using a retrieval-based few-shot in-context learning method.
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4 & §6: The Amount of Mod Discourse for a given community is the fraction of all posts and comments discussing mods in that community, regardless of their sentiment. The Composition of Mod Discourse with positive/negative sentiment is the percent of all posts and comments discussing mods in a given community that have positive/negative sentiment. To ensure a meaningful amount of data, we exclude communities that did not have at least one positive, neutral, and negative post or comment discussing moderators. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. This filtering leaves us with 5,282 subreddits used in downstream analyses.
4. How are moderators of different communities perceived differently by their communities?
Online communities exist for nearly every conceivable topic, and range in size from just a few members to many millions. In this section, we examine how community members’ perceptions of their moderators vary across communities with different topics, different community health metrics, and different sizes.
Method. To quantify a community’s health, we divide communities into quartiles based on their community members’ responses to a recent survey (§3.2). Survey responses were collected between May-July 2021, a period overlapping the the end of our mod discourse data time range. For this analysis only, we exclude communities which were not surveyed. Community topic is classified into one of six topical categories using our topic classifier (§3.2). We quantify community size as the volume of submitted content: the average number of posts and comments per day over our study period, which we use to group communities with similar size. For each group, we compute the amount of mod discourse as well as the composition of that discourse.
Results. Different aspects of community health are associated with better and worse perceptions of moderators. The smallest-feeling 25% of communities, based on member self-reports, use 27% more positive and 16% less negative sentiment to discuss their moderators than the largest-feeling 25% of communities (Figure 2a). Figure 2a-e shows that communities which rate themselves as having higher quality content and being more trustworthy, more inclusive, and more safe all use more positive and less negative to discuss their moderators, as well.
Figure 3. Perceptions of moderators vary significantly across communities with different sizes (a-b) and topics (c-d). In general, smaller communities devote a relatively larger proportion of their content to discussing their moderators (a), and smaller communities express more positive and less negative sentiment towards their mods (b). Discussion, meme-sharing, and news communities have proportionally more mod discourse (c), while meme and news-sharing communities exhibit the most negative sentiment towards their moderators (d).
Direct measurement of communities’ size allows us to further investigate the relationship between the volume of content submitted to a community, and its members’ perceptions of its moderators. In general, smaller communities have more mod discourse (Figure 3a), with tiny communities with $<10$ posts+comments/day having 7.6×7.6$\times$7.6 × as many posts and comments discussing moderators as huge communities (those with more than 10k posts and comments per day), relative to the total amount of content. Tiny communities also have 6.0×6.0$\times$6.0 × more positive mod discourse, and 0.46×0.46$\times$0.46 × as much negative discourse, than huge communities.
Examining communities of different topics, discussion communities have the most mod discourse, on average (Figure 3c). Hobby communities have generally the most positive mod discourse, with meme and news-sharing communities having 46% more of their mod discourse have negative sentiment than communities of other topics (Figure 3d).
Implications. Surveys of redditors have shown that they consider a wide range of factors to be important to the overall ‘health’ of their communities (Weld et al., 2024). Here we show that communities that are doing well with regards to factors widely considered to be important, such as safety and quality of content, tend to perceive their moderators more positively. Many of these factors are only indirectly controlled by the moderators, for example quality of content—while moderators can set and enforce rules aimed at improve the quality of content in their communities, quality of content is ultimately a function of the content submitted by community members, not moderators. This can lead to moderators being ‘blamed’ for problems largely outside of their control, which has been previously shown in small-n surveys (Matias, 2019a; Jhaver et al., 2023a) and is further supported empirically here, as shown in Figure 2.
Small communities appear to both discuss their moderators more, as well as use more positive sentiment in their discussions (Figure 3a,b). Several factors may contribute to this, including that the increased anonymity that comes with participating in a larger community may make people feel more comfortable speaking negatively about the moderators, and that smaller communities are more likely to be newly formed and thus still establishing moderation norms, leading to more mod discourse (Hwang and Foote, 2021; Seering and Kairam, 2022).
5. Adjusting for Confounders using IPTW & DID
While we are ideally interested in causal, actionable insights, our study is a retrospective observational study, and like any other, subject to potential confounding. In §4 we identified two important confounders, size and topic. Next we describe two methods from the causal inference literature which we use to adjust for these and other confounding factors in subsequent analyses.
Inverse Probability of Treatment Weighting (IPTW). IPTW is a statistical method to adjust for imbalance in possible confounding factors when making comparisons between different observed groups (e.g., a treatment and a control group). We select IPTW over other methods, such as stratification, due to its excellent efficiency (Austin and Stuart, 2015). In §6, we compute the probability of treatment (propensity score, $P(Z|\mathbf{X})$) by applying a logistic regression model to community covariates $\mathbf{X}$ including the topic and size of the community. A complete list of covariates for each analysis is given in Supplementary Materials  C. We generalize IPTW to non-binary treatments to estimate a dose-response curve using the method from (Althoff et al., 2016) in which the treatment is discretized into bins and a different propensity score is computed for each treatment bin in a ‘one-versus-rest’ scheme. For final analyses, each observation is weighted by the inverse of the probability of the treatment it received, such that the weight for observation $i$ is $w_{i}=\frac{Z_{i}}{P(Z_{i}=1|\mathbf{X}_{i})}+\frac{1-Z_{i}}{P(Z_{i}=0|\mathbf{X}_{i})}$, where $Z_{i}$ is the treatment received by observation $i$ (0 for control, 1 for treated), and $\mathbf{X}_{i}$ is a vector of the covariates. In analyses where we use IPTW, we also include the non-IPTW adjusted estimate in gray. In these analyses, reweighting does not dramatically impact our findings, suggesting that the impact of these confounders is moderate at best.
An important validity check for IPTW is to assess the balance of covariates for each treatment group after weighting as applied (Austin and Stuart, 2015). Two groups are often considered ‘balanced’ or ‘indistinguishable‘ if all covariates are within a standardized mean difference (SMD) of 0.25 standard deviations (Althoff et al., 2016). For each treatment group, we compute the difference between each covariate’s weighted mean value and the reference distribution, consisting of the entire population. We compute the standardized mean difference (SMD) by normalizing the difference in means by the standard deviation of the values of the reference distribution. Our reweighting method achieves balance for every covariate substantially related to the treatment. SMDs after weighting are given for each covariate and each analysis in Supplementary Materials C.
Difference in Difference Analyses (DID). Difference in difference analyses enable the estimation of the impact of an intervention by comparing the outcome before vs. after the intervention was applied to the treated group, and comparing this before vs. after difference to an untreated control group. For our analyses of the impact of moderators’ community engagement (§6.2), we use modified DID analyses with multiple time periods (Callaway and Sant’Anna, 2021) to compare the difference in communities’ receptions of new moderators who are engaged with the community (the ‘treatment’) vs. those who are not (the ‘control’). Two four week long periods immediately preceding and following each moderators’ appointment were used to compute the first-order difference. Since mod discourse naturally varies over time, averaging over many DIDs from many new moderator appointments serves to reduce confounding background temporal trends by comparing the difference to an untreated control group.
6. What can moderators do to improve how they are perceived by their community?
Moderators have a great deal of autonomy to run their communities as they see fit, including enforcing rules by removing content and growing the mod team by recruiting or appointing new moderators. Moderators also may (or may not) participate in the community as ‘regular community members’ in addition to their mod duties. In this section, we identify promising suggestions for moderators by comparing communities whose moderators have different workloads, rule enforcement strategies, and degrees of community engagement.
6.1. Content Removal and Moderator Workload
Method. We can identify content removed by moderators in each community by counting the occurrences of ‘[removed]’ posts and comments within each community. By dividing by the total amount of content submitted to that community, we can compute the total percentage of content removed by mods. Using our mod timelines (§3.1), we can compute the total number of moderator-tenures in any given subreddit over the course of our analysis time period. We also approximate ‘workload’ of each mod by dividing the total amount of content submitted to each community by the number of mods available to review that content; while in reality it is unlikely that all moderators share the work of reviewing content evenly (Li et al., 2022a); this metric nonetheless helps us understand the ratio of moderators to content within each community. If workload was shared unevenly, this would only make the workload even higher for a subset of moderators, leading to at worst to overly conservative estimates (higher workload to be addressed by fewer moderators).
Results. We find that communities with higher mod workloads have more negative mod discourse (Figure 4). Communities with fewer than five pieces of daily content for each mod (8.8% of all communities) have approximately equal amounts of mod discourse with positive and negative sentiment, a rarity on a platform where mods are far more commonly discussed negatively. In communities with higher mod workload, the composition of mod discourse is much less positive, with communities with more than 100 posts and comments per mod per day using positive sentiment to describe their mods only 0.39×0.39$\times$0.39 × as often as the 8.8% of communities with the lowest mod workload.
Figure 4. Moderators in communities with lower workloads are perceived more positively and less negatively than moderators in communities with high workloads. Communities with lower moderator workloads (more moderators relative to the amount of content submitted) tend to have more more positive sentiment in their discussion of the moderators, and less negative sentiment. Communities with fewer than five posts and comments per mod per day use 2.5×2.5$\times$2.5 × as much positive sentiment in their mod discourse compared to communities with more than 100 posts and comments per mod per day.
Generally, we find that the composition of mod discourse is more negative in communities with more removed content (Supplementary Figure 9). However, these trends vary depending on the topic of the community, with news communities in particular showing the opposite trend (Figure 5). [^2] Similar figures for each of the six community topics are included in Supplementary Figure 10. Amongst news communities that remove $\leq 1\%$ of their content, 48% of mod discourse has negative sentiment, which drops 11pp to 39% for news communities whose mods remove between 2% and 3% of their content. For the same amounts of removed content for non-news communities, the fraction of mod discourse with negative sentiment increases 6pp, from 26% to 33%.
Figure 5. For most topics, communities where moderators remove more content exhibit more negative sentiment (a). News communities, however, buck this trend, with the fraction of mod discourse with negative sentiment 11pp lower in news communities whose mods remove less than $1\%$ of content compared to communities whose mods remove 2% - 3% of content (b).
Implications. Our results suggest that adding additional moderators to a community (and therefore reducing the effective moderator workload) may improve community members’ perceptions of their mods (Figure 4). In §6.2 we report on additional longitudinal evidence that recruiting additional moderators can have a positive effect. However, simply increasing the amount of content which is removed does not, in general, appear to be associated with more positive moderator discourse sentiment—in fact, the opposite appears to be the case for communities that are not focused on sharing news (Figure 5). Taken together, these results imply that there are topic-specific nuances to content removal, and that mods should use care when deciding how strictly to enforce rules, and how much content to remove. Our results suggest that certain topics are more amenable to stricter rule enforcement than others.
6.2. Community Engagement
Method. Some moderators are actively engaged with community, soliciting feedback from non-moderators, updating community members on moderation-related news, and contributing to regular community content in addition to their official moderator duties. Other moderators are far less visible, opting to remove content and change rules without participating in the community more broadly. Furthermore, many moderators also participate in other communities beyond just the one(s) they moderate. Using moderators’ public posts and comments, for each moderator-appointment to a community, we compute the number of posts and comments they made in that community and in other communities before and during their tenure as a moderator. We use DID analyses (§5) to estimate the impact of appointing new mods with different degrees of community engagement (or lack thereof).
Results. We find that, regardless of the moderators’ previous engagement with the community, adding a new moderator to a community is associated with a decrease in mod discourse with negative sentiment (Figure 6). However, not all moderators have the same impact. The most positively impactful moderators are those who are engaged with the community prior to their appointment (Figure 6a), will continue to engage with the community during their tenures (Figure 6b), and were also active in other communities prior to their appointment (Figure 6c). New mods who do all three of these things are associated with a 2.5pp increase in positive mod discourse, whereas mods who do none of these things are associated with a 0.5pp decrease.
Implications. Our results suggest that moderators who actively engage with the community both before and during their tenures are most likely to have a positive impact on the community’s perception of their mod team, which may be because these moderators are more familiar with the community, and are better mods as a result. Another plausible mechanism for this effect is that non-moderator community members value the transparency and accountability that may stem from increased moderator engagement. Our results also suggest that moderators who are active in other subreddits beyond the one(s) they moderate have a more positive impact on the communities they moderate, perhaps because participating in a broader range of communities makes them more effective moderators (Zhu et al., 2014). Lastly, our results suggest that of these factors, engagement with the community during the moderator’s tenure has the largest impact.
Figure 6. Newly appointed mods are associated with a greater improvement in mod perceptions if they are engaged in the community and elsewhere on reddit before their tenure, and if they are engaged during their tenure (a-c). Adding a moderator who already has or will engage with the community is associated with in an increase in the fraction of mod discourse in the community with positive sentiment (aqua), and that increase is 32.5% larger when adding a mod who will engage with the community going forward (b) than for one who already has (a). Adding a moderator who is an active member of communities other than the one they are becoming a mod of is associated with an increase in positive, and a decrease in negative, sentiment in mod discourse (c).
6.3. Moderator Recruiting
When it’s time to grow the mod team, existing moderators have a wide array of options for who to recruit, and how to recruit them, and little guidance, official or otherwise, for how to select new mods. In practice, moderator recruiting falls into two different strategies: Public Recruiting, where moderators post publicly internally in their own subreddit that they are looking for moderators, and solicit applications, nominations, or hold elections. Sometimes, moderators make use of special external moderator-recruiting subreddits, such as /r/needamod, where they can post ‘job listings’ to prospective applicants. By contrast, Private Recruiting is the recruiting of new moderators in the absence of any public recruiting activity. Moderators recruited privately are either invited to join directly by an existing mod, make an offer to moderate themselves, or are recruited via other backchannel means.
Figure 7. Communities that recruit moderators publicly are 8.78×8.78$\times$8.78 × larger than the average community which recruits only privately, in terms of the community’s daily volume of content (a). Small communities lean towards private recruiting. (b) Compared to private recruiting, recruiting moderators publicly is polarizing: it is associated with an increase in both positive and negative fractions of mod discourse, and a corresponding decrease in neutral sentiment.
Method. We identify instances of public recruiting using a regular expression-based search for posts made by sitting moderators which use ‘recruiting new mods,’ ‘applications open for new mods,’ ‘holding mod elections,’ or similar phrases. We also apply a regular expression-based filter to the complete set of posts from /r/needamod to identify which subreddits are recruiting, restricting the results to only posts made by current moderators of the subreddit that is recruiting. We then take our dataset of moderator timelines (§3.1) and match moderators who were appointed to a community within 8 weeks of a public recruiting post as having been publicly recruited. Moderators appointed without a recent public recruiting post are considered privately recruited.
Results. We find that, while communities of all sizes use public recruiting methods, larger communities are the most likely to make use of public recruiting (Figure 7a). In terms of its daily volume of content, the average community that uses both Internal and External Public Recruiting is 8.78×8.78$\times$8.78 × larger than the average community that only recruits privately. Examining the impact that adding a single moderator recruited privately vs. publicly has to a community in Figure 7b, we find that public recruiting appears to be polarizing, with the fractions of both positive and negative mod discourse increasing by 2.7pp and 3.0pp, respectively, after a publicly recruited mod is added, on average.
Implications. Our findings suggest that public moderator recruiting has the potential be a powerful tool to improve community perceptions of moderators when used carefully, and can be harmful when used without regard to the community’s preferences. A plausible explanation for the increased polarization resulting from public recruiting is that in some circumstances, the public nature of the recruiting exacerbates existing frustrations with mod teams, for example if due process was not followed during moderator elections, or if a perceived-outside was brought in via external recruiting when community members themselves preferred someone with more experience in the community. More work is needed to assess the differences between different moderator recruiting strategies.
6.4. Novice and Experienced Mods
Method. When selecting new moderators, existing moderators may be inclined to favor candidates who already have some moderation experience in other communities. Using Moderator Timelines (§3.1), we can accurately assess how much experience a new moderator has at the time the are appointed, if any.
Figure 8. In contrast to small moderator teams, large teams are much more likely to appoint new moderators who already have moderation experience. $94\%$ of mods recruited to join mod teams with more than 100 mods have at least 2 years of experience, while 74% of mods who join small teams with fewer than 10 mods have no previous experience at all.
Results. We find that large communities are much more likely to recruit mods who already have moderation experience (Figure 8). $74\%$ of mods appointed to communities with fewer than 10 mods are first time mods, while $94\%$ of mods appointed to communities with $>100$ mods have more than 2 years of experience.
Implications. Large communities with large moderator teams rarely appoint novice moderators, perhaps due to a perception that a large community is not an appropriate place for new moderators to gain experience. As such, it seems that the most common pattern on reddit is for new moderators to start off moderating small communities, and then work their way up to larger ones. This may contribute to a perception, especially amongst non-moderators, that moderators are motivated to increase their number of appointments and to moderate larger and larger communities, potentially biasing the performance of their duties.
7. Discussion
The massive diversity of online communities offers enormous potential to empirically study how to make online communities better. Any such studies, however, require robust and scalable methods to quantify both the outcomes (which communities are doing ‘better’ than others), as well as the independent variables (the aspects of communities that might make them better). Measuring the success of a community’s moderators is more challenging than many other aspects of making a community ‘better,’ as unlike, say, misinformation, governance is far less visible in a community, and its success is far less well defined. The primary contribution of our work is the use of community members’ own perceptions of their moderators, leveraging millions of people’s perceptions of good governance, rather than attempting to define good governance ourselves.
We are tremendously excited about the potential synergies this line of work enables. While the findings presented in this paper are impactful by themselves, combining our measure of governance with many other important measures of community outcomes (safety, inclusion, and discussion quality, for example) will enable studies of how to make communities better that are both more comprehensive and more robust than previously possible.
7.1. Limitations
Our work measures community governance through community members’ public discussion of moderators. While these signals enable insights about different governance strategies, they do not capture every aspect of the success of a community’s governance. What people say publicly does not always reflect their actual beliefs, and even if it did, minimizing community unhappiness is not necessarily the best objective function for community moderators. Differences in community members’ behavior and values also may bias our results, as different community may have different norms around mod discourse, and different community members may feel more or less comfortable expressing their opinions publicly. Additional study is needed to ensure that scalable methods for measuring governance reflect the needs of all community members, not just the noisiest.
Even though our fine-tuned sentiment classifier exceeds the performance of state-of-the-art methods (§3.3), it is not perfect. The large volume of data of data used in our analyses minimizes the impact of any single misclassification by the model, and all of our figures include bootstrapped CIs to better understand the robustness of our findings. Furthermore, systems as complex as online communities have countless confounding factors that can bias analyses such as ours. While we attempt to control for confounders by using IPTW and DID causal inference methods (§5), future work should make use of active experimentation for gold-standard causal estimates.
Validating measures of online governance is challenging, as there is minimal ‘ground truth’ to use for assessment. While it exceeds the scope of this work, a large-scale survey of many community members’ perceptions of their moderators could be used to refine and validate future models. Future models could go beyond just positive and negative sentiment to identity biased, overly-strict, or too-permissive moderation, for example.
While our methods for quantifying communities’ practices are sophisticated, they also miss many key aspects of governance. Our analyses of moderators’ engagement with communities (§6.2) only consider how much each moderator posts in their communities, not what they post. Future analyses could examine the types of contributions that moderators make. Our analyses of moderator team dynamics are also limited by the lack of publicly available data about which moderators are active; many communities’ mod teams contain mods who do not contribute to the day-to-day governance of the community. Future work should incorporate detailed information about specific moderators’ actions, although data collection is a substantial challenge (Li et al., 2022a).
7.2. Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across thousands of communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§6.1), community engagement (§6.2), and moderator recruiting practices (§6.3-6.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our models and datasets public to support future research.
8. Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes, 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. This study was approved by the University of Washington IRB under ID number STUDY00011457.
Supplementary Materials
Appendix A Moderator Sentiment Codebook
A.1. Positive Sentiment
This label should be used for comments expressing a positive sentiment towards the moderator or moderator team.
Examples. “This subreddit is so lucky to have such a great mod team”
“Make the life of our hard-working mods here easier”
“The mods are always so helpful, but this thread got a bit messy” — this is a tricky judgment call, but I’d say that the overall sentiment is positive with this thread being an exception.
Counterexample. “This subreddit used to be well-run, but in the past year or so the moderation has really gone to shit” — a judgment call similar to earlier, but here I would say negative.
A.2. Negative Sentiment
This label should be used for comments expressing a negative sentiment towards the moderator or moderator team.
Examples. “The mods here suck”
“The mods made a mistake” — everyone makes mistakes, but it’s still better if they don’t.
“I’m so tired of mods not removing crap like this”
A.3. Neutral Sentiment
This label should be used when there isn’t enough context for you to make a judgment about the sentiment of the comment or post, or the sentiment seems neutral.
Examples. “I didn’t delete the post, maybe the mods did?”
“Edit: reworded a slur after getting a warning from the mods”
“Mods please ban this person” — not enough explicitly stated sentiment to know what is meant with certainty.
“Why don’t you go and complain to the mods like you usually do?” — negative sentiment, but not directed towards the moderators.
Appendix B Supplementary Figures
Figure 9. On average, communities with a larger fraction of their content removed by mods tend to have more a smaller fraction of their mod discourse have positive sentiment (b), and a larger fraction with negative sentiment (c). Communities with more content removed by mods also tend to have more total mod discourse (a).
Figure 10. Generally, the fraction of mod discourse with negative sentiment is higher in communities with more removed content. However, these trends vary depending on the topic of the community.
Figure 11. On average, adding a new moderator to a subreddit results in an increase in the fraction of mod discourse which has positive sentiment, and a corresponding decrease in the fraction that has negative sentiment. However, the magnitude of the impact varies with the size of the community’s moderator team; adding a single new mod to a community with fewer than 4 mods has a much larger impact than adding a single mod to a community with more mods. Adding a single moderator to a community with more than 15 mods has an impact which is not significantly different from 0.
Appendix C Covariates Used in Propensity Score Modeling for IPTW
C.1. Covariates used for Mod Workload Analyses
The following table shows the covariates that were used in a logistic regression propensity score model for the moderator workload analyses (Figure 4), along with the resulting Standardized Mean Differences (SMDs) after reweighting, and the associated propensity score model (P.S.) coefficients used in the logistic regression.
Moderator Workload Treatment Bin (Posts+Comments per Mod per Day)
0-5 5-10 10-100 ¿100
SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient
Covariate
total_items -0.18 -0.55 -0.17 -0.45 -0.10 +0.07 +0.48 +0.93
frac_deleted +0.05 +0.01 +0.02 +0.02 -0.07 -0.07 +0.20 +0.04
frac_removed +0.19 +0.04 +0.07 +0.01 -0.06 -0.05 -0.03 +0.00
num_mods +0.04 +0.29 -0.02 +0.23 -0.03 -0.08 -0.01 -0.44
category_hobby -0.22 -0.02 +0.01 +0.01 +0.06 +0.02 -0.07 -0.01
category_discussion +0.18 +0.01 -0.02 -0.01 -0.05 -0.01 +0.04 +0.01
category_memes +0.18 +0.03 +0.01 +0.00 -0.04 -0.05 +0.00 +0.01
category_news +0.03 -0.01 +0.01 +0.00 -0.01 +0.01 +0.00 -0.01
category_media +0.34 +0.02 +0.10 +0.00 -0.08 -0.04 -0.00 +0.01
category_identity -0.24 -0.04 -0.10 -0.01 +0.05 +0.07 +0.07 -0.02
C.2. Covariates used for Strictness of Rule Enforcement Analyses
The following table shows the covariates that were used in a logistic regression propensity score model for the rule enforcement analyses (Figure 5), along with the resulting Standardized Mean Differences (SMDs) after reweighting, and the associated propensity score model (P.S.) coefficients used in the logistic regression.
Amount of Removed Content Treatment Bin (Percentage of All Content)
0%-1% 1%-2% 2%-3% ¿3%
SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient
Covariate
total_items -0.10 -0.12 +0.03 -0.10 +0.12 +0.01 +0.12 +0.21
frac_deleted -0.58 -0.51 +0.02 -0.01 +0.30 +0.10 +0.67 +0.42
mod_workload -0.08 -0.00 +0.04 +0.14 +0.15 +0.02 +0.04 -0.16
num_mods -0.06 -0.21 -0.01 +0.12 -0.01 +0.01 +0.16 +0.08
category_hobby +0.53 +0.10 -0.17 -0.01 -0.29 -0.01 -0.46 -0.08
category_discussion -0.26 -0.09 +0.00 +0.01 +0.02 -0.01 +0.40 +0.09
category_memes -0.04 +0.08 -0.01 -0.00 +0.04 +0.01 -0.02 -0.08
category_news -0.11 -0.12 -0.06 -0.06 -0.00 -0.00 +0.28 +0.19
category_media -0.12 +0.03 -0.08 -0.01 +0.02 -0.00 +0.21 -0.02
category_identity -0.33 +0.00 +0.32 +0.08 +0.31 +0.02 +0.02 -0.10